# Choose Wisely: Data-Enabled Predictive Control for Nonlinear Systems Using Online Data Selection

### v1

@@ 5 Conclusion
This paper introduces Select-DeePC, a novel approach to constraint-aware Data-Driven Predictive Control for nonlinear systems. At each time instance, the method pre-processes the data set to determine the most relevant data, then passes the most relevant data to a convex optimization which determines the optimal control for the future horizon. The optimal control for the given timestep is implemented, and the algorithm progresses in a receding horizon fashion.
In contrast to other DDPC methods, our method can be implemented easily, scales favorably with the number of collected data-points and increases performance outside of the linear domain.
We validated this in a variety of forms and on a set of three simulation environments. In the rocket simulator environment, we showed that for one of the data-sets considered, Select-DeePC cut closed-loop cost by a factor of 2. For the other data-set, our method succeeded, where previous methods crashed. In the second set of simulations in the Reacher environment, we enabled zero-shot constrained setpoint tracking of a reference signal in a robotic reacher simulation. Again, this was not achievable using previous DeePC formulations. Finally, in the last set of simulations of a cart-pole inverted pendulum, we demonstrated successful transitioning between two different equilibria of the system by performing a pendulum swing-up.
A direction of future work is to investigate different selection methods, such as structure-exploiting selection methods or selection methods based on an information criterion. Another direction is to replace Select-DeePC’s large training data set with a data-driven forward simulator of the dynamics (i.e., a “world model”), resulting in a constraint-aware alternative to Sampling-Based MPC.
Acknowledgments
The authors would like to thank Xian Li for providing the simulation environment and data collection procedure for the inverted pendulum.
@@ Appendix A DeePC
Lemma A.2 (The Fundamental Lemma (Willems et al., 2005)).
@@
Here, $r(g)$ denotes suitable regularizers on $g$ (Dörfler et al., 2023) which improve robustness of the controller in the presence of noise. The implicit predictor in (5) can readily be extended to affine systems as follows (Martinelli et al., 2022; Berberich et al., 2022)
@@ Appendix B SQP-MPC
The full SQP algorithm can be summarized as follows
@@
B.1 Equivalence between Select-DeePC and SQP-MPC
Select-DeePC has a very natural interpretation as solving a nonlinear MPC problem using SQP, where the analytical Jacobian computation is replaced by a data-driven approach of estimating the Jacobian of the input-output behavior from data. Specifically, one can select data points “close” to the current linearitation point (given by the open loop solution), resulting in an equivalent, data-driven approximation of $F_{\text{p}}$ and $F_{\text{f}}$ in Equation 14 that approaches the analytical Jacobians assuming sufficient amounts of data in the neighbourhood around the linearization point (i.e. $\operatorname{card}(\mathcal{D})$ tending to infinity).
@@ Appendix C Isomap
Isomap, short for Isometric Mapping, is a nonlinear dimensionality reduction technique introduced by Tenenbaum (2000). It can be seen as an extension of classical Multidimensional Scaling (MDS) where, instead of using Euclidean interpoint distances, geodesic distances between the data points are preserved. These geodesic distances are estimated using shortest distances on a connected graph which connects neighboring data points. These shortest paths are then arranged in a distance matrix $D$. Isomap then finds an embedding by applying classical MDS. Computing the distance of a new query data point to the data set in the embedding dimension is done by first linking the new data point into the neighborhood graph and then projecting the data point into the embedding space.
Isomap is able to capture the overall nonlinear nature of the data manifold and is guaranteed to produce a global optimizer of its reconstruction loss. Furthermore, due to preserving geodesic distances, points close in the input space are also close in the embedding space, resulting in interpretable results. This is especially relevant for Select-DeePC which is build upon the idea of selecting data points according to their associated distance to a query point.
@@ C.1 Isomap Parameter Selection
Furthermore, note the sharp increase in reconstruction error as the number of graph neighbors increases after the local minimum. This demonstrates the effect of bad hyperparameter selection which can lead to short-circuiting or over-connectedness in the graph, leading to worse embedding performance as the local manifold structure is lost (Tenenbaum et al., 2000).
@@ Appendix D Experiment Hyperparameters
where $Q$ and $R$ are positive (semi-)definite cost matrices and $y_{r}$ is a reference setpoint. Furthermore, for the regularization term $r(g)$ we use 1-norm and projection regularization (Markovsky et al., 2023) with weights $\lambda_{1}$ and $\lambda_{\Pi}$ respectively. The following tables shows the hyperparameters for the three sets of simulations.

### v2

@@ 5 Conclusion
This paper introduces Select-DPC, a novel approach to constraint-aware Data-driven Predictive Control for nonlinear systems. At each time instance, the method pre-processes the data set to determine the most relevant data, then passes the most relevant data to a convex optimization which determines the optimal control for the future horizon. The optimal control for the given timestep is implemented, and the algorithm progresses in a receding horizon fashion.
In contrast to other DPC methods, Select-DPC can be implemented easily, scales favorably with the number of collected data-points and increases performance outside of the linear domain. Furthermore, compared to Sampling-Based MPC, Select-DPC can explicitly handle output constraints.
We validated this in a variety of forms and on a set of three simulation environments. In the rocket simulator environment, we showed that for one of the data sets considered, Select-DPC cut closed-loop cost by a factor of 2. For the other data set, our method succeeded, where previous methods crashed. In the second set of simulations in the Reacher environment, we enabled zero-shot constrained setpoint tracking of a reference signal in a robotic reacher simulation. Again, this was not achievable using previous DPC formulations. Finally, in the last set of simulations of a cart-pole inverted pendulum, we demonstrated successful transitioning between two different equilibria of the system by performing a pendulum swing-up.
A direction of future work is to investigate different selection methods, such as structure-exploiting selection methods or selection methods based on an information criterion. Another direction is to replace Select-DPC’s large training data set with a data-driven forward simulator of the dynamics (i.e., a “world model”), resulting in a constraint-aware alternative to Sampling-Based MPC.
@@ Appendix A DeePC
Lemma A.2 (The Fundamental Lemma Willems et al. (2005)).
@@
Here, $r(g)$ denotes suitable regularizers on $g$ Dörfler et al. (2023) which improve robustness of the controller in the presence of noise. The implicit predictor in (5) can readily be extended to affine systems as follows Martinelli et al. (2022); Berberich et al. (2022)
@@ Appendix B SQP-MPC
The full SQP algorithm can be summarized as follows Diehl et al. (2002)
@@
B.1 Equivalence between Select-DPC and SQP-MPC
Select-DPC has a very natural interpretation as solving a nonlinear MPC problem using SQP, where the analytical Jacobian computation is replaced by a data-driven approach of estimating the Jacobian of the input-output behavior from data. Specifically, one can select data points “close” to the current linearitation point (given by the open loop solution), resulting in an equivalent, data-driven approximation of $F_{\text{p}}$ and $F_{\text{f}}$ in Equation 14 that approaches the analytical Jacobians assuming sufficient amounts of data in the neighbourhood around the linearization point (i.e. $\operatorname{card}(\mathcal{D})$ tending to infinity).
@@ Appendix C Isomap
Isomap, short for Isometric Mapping, is a nonlinear dimensionality reduction technique introduced by Tenenbaum Tenenbaum et al. (2000). It can be seen as an extension of classical Multidimensional Scaling (MDS) where, instead of using Euclidean interpoint distances, geodesic distances between the data points are preserved. These geodesic distances are estimated using shortest distances on a connected graph which connects neighboring data points. These shortest paths are then arranged in a distance matrix $D$. Isomap then finds an embedding by applying classical MDS. Computing the distance of a new query data point to the data set in the embedding dimension is done by first linking the new data point into the neighborhood graph and then projecting the data point into the embedding space.
Isomap is able to capture the overall nonlinear nature of the data manifold and is guaranteed to produce a global optimizer of its reconstruction loss. Furthermore, due to preserving geodesic distances, points close in the input space are also close in the embedding space, resulting in interpretable results. This is especially relevant for Select-DPC which is build upon the idea of selecting data points according to their associated distance to a query point.
Figure 7: Relative contrast with Isomap embedding compared to $L_{1}$ selection.
The relative contrast of a data set ${\Delta=(d_{\text{max}}-d_{\text{min}})/d_{\text{min}}}$, with $d_{\text{max/min}}$ indicating the maximum and minimum norm values of the data set respectively, is a helpful surrogate to determine how distinct each data point appears under a given distance metric. Figure 7 shows the relative contrast of a data set collected from the rocket simulator in Section 4 as a function of trajectory dimensionality when embedded using Isomap. It compares relative contrast given a fixed embedding dimensionality as well as a varying one, which increases according to $n+T_{\text{f}}\cdot m$, to the relative contrast of the $L_{1}$ norm. Both Isomap methods have better relative contrast compared to $L_{1}$ selection. While a varying embedding dimensionality offers greater relative contrast for small trajectory lengths, we encounter diminishing returns for large trajectory lengths.
@@ C.1 Isomap Parameter Selection
Furthermore, note the sharp increase in reconstruction error as the number of graph neighbors increases after the local minimum. This demonstrates the effect of bad hyperparameter selection which can lead to short-circuiting or over-connectedness in the graph, leading to worse embedding performance as the local manifold structure is lost Tenenbaum et al. (2000).
@@ Appendix D Experiment Hyperparameters
where $Q$ and $R$ are positive (semi-)definite cost matrices and $y_{r}$ is a reference setpoint. Furthermore, for the regularization term $r(g)$ we use 1-norm and projection regularization Markovsky et al. (2023) with weights $\lambda_{1}$ and $\lambda_{\Pi}$ respectively. The following tables shows the hyperparameters for the three sets of simulations.